OpenStudy (anonymous):

(7,1) and perpendicular to y= -x+3 help!!!!!!!!

OpenStudy (anonymous):

If u want a line perpendicular to another line help this : m1.m2=-1 Means m2=1/x Got it ?:)

OpenStudy (anonymous):

using\[y = mx + b\] we can determine the slope of the original equation which is \[m = -1\] to find the equation of the line perpendicular to the original equation passing through (7,1), we will need the negative reciprocal of determined slope. In this case, \[m _{p} = -(1/(-1)) = 1\] now we have a point and a slope so we need to use the point slope form \[(y-y _{1}) = m _{p}(x-x _{1})\] solve for y to get your solution....let me know if you need further assistance.